Luzadis Company makes furniture using the latest automated technology. The company uses a job-order costing system and applies manufacturing overhead cost to products on the basis of machine-hours. The predetermined overhead rate was based on a cost formula that estimates $592,000 of total manufacturing overhead for an estimated activity level of 74,000 machine-hours.

During the year, a large quantity of furniture on the market resulted in cutting back production and a buildup of furniture in the company’s warehouse. The company’s cost records revealed the following actual cost and operating data for the year:

Machine-hours 67,000
Manufacturing overhead cost $ 551,000
Inventories at year-end:
Raw materials $ 13,000
Work in process (includes overhead applied of $37,520) $ 139,300
Finished goods (includes overhead applied of $101,840) $ 378,100
Cost of goods sold (includes overhead applied of $396,640) $ 1,472,600

Required:

1. Compute the underapplied or overapplied overhead.

2. Assume that the company closes any underapplied or overapplied overhead to Cost of Goods Sold. Prepare the appropriate journal entry.

3. Assume that the company allocates any underapplied or overapplied overhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry.

4. How much higher or lower will net operating income be if the underapplied or overapplied overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather than being closed to Cost of Goods Sold?
Answer #1
Actual machine hours 67000
X Predetermined overhead rate 8 =592000/74000
Overhead applied 536000
1
Manufacturing overhead cost 551000
Less: Overhead applied 536000
Underapplied overhead 15000
2
General Journal Debit Credit
Cost of goods sold 15000
        Manufacturing overhead 15000
3
General Journal Debit Credit
Work in Process 1050 =15000/536000*37520
Finished goods 2850 =15000/536000*101840
Cost of goods sold 11100 =15000/536000*396640
        Manufacturing overhead 15000
4
Net operating income will be $3900 greater if the underapplied overhead (15000-11100)
is allocated rather than closed entirely to cost of goods sold